Equation of hyperbola with asymptotes are given as 3x-4y+7=0 and 4x+3y+1=0?

Grade:12

1 Answers

Arun
5 years ago
 

Hence, the joint equation of the asymptotes is (3x – 4y + 7)(4x + 3y + 1) = 0.

Now, we know that the equation of rectangular hyperbola differs from that of the joint equation of asymptotes by a constant only, hence we must have the equation of the hyperbola as

(3x – 4y + 7)(4x + 3y + 1) + r = 0.

It is given that the hyperbola passes through the origin and hence, putting x = y = 0 , we obtain,

7+ r = 0 

Hence, r = -7.

Hence, the equation of the hyperbola is (3x – 4y + 7)(4x + 3y + 1) – 7= 0.

This gives 12x2 + 9xy + 3x -16xy – 12y2 – 4y + 28x + 21y + 7- 7 = 0.

So, 12x2 – 12y- 7xy + 31x + 17y = 0